Several monochromatic lights pass through double slits with a separation of s 50 3m. The screen to slits distance is 2.5 m. What is/are the width of its central maximum and its corresponding wavelength(s) where the bright spot can occur? [width is expressed in 3 significant figures]

Wavelength (nm) Width (cm)

A. 350 1.75

B. 400 2.00

C. 575 5.75

D. 700 3.50


1
Expert's answer
2021-07-19T15:43:21-0400

Gives

"d=503nm \\D=2.5m\\\theta=53°"

Central bright fringe

"m\lambda=dsin\theta"

Put value

m=1

"\lambda=503\times sin53°"

"\lambda=400.17" nm

"\lambda=400nm"

Fring width

"\beta=\frac{\lambda D}{d}=\frac{400\times10^{-9}\times2.5}{503\times10^{-9}}=2.0"


Option (b) is correct option